# GraphEx: A Graph-based Extraction Method for Advertiser Keyphrase Recommendation

### v2

@@ 3.3. Terms and Notations
We consider a set of unique keyphrases termed as $Q={k_{1},k_{2},...,k_{K}}$. Each keyphrase $k_{i}$ can be considered as a set of words ${w_{1},w_{2},...,w_{l}}$, where $w_{*}$ are tokenized [^4] The tokenization scheme can be anything as long as string comparison functions are well-defined and consistent for that scheme. By default we consider space delimited tokenization. from the keyphrase string $k_{i}$. Each $k_{i}$ is further associated with a Leaf category $l$, Recall Count/Rank $R$ and Search Count/Rank $S$. Given a test item’s title $T$, the goal is to recommend a subset of keyphrases from $Q$ that are relevant to $T$. We can consider the title as a string with tokenized4 words $T={w_{1},w_{2},...,w_{t}}$ similar to a keyphrase, but titles are generally longer than the keyphrases. We denote a graph $G(V,E)$ where $V$ is the set of vertices and $E$ is the set of edges. Each edge $e\in E$ is denoted by a pair of vertices $e=(v_{1},v_{2})$ indicating a connection between the vertex pair. In a Bipartite Graph, the set of vertices $V$ are divided into a pair of disjoint subsets $V=X\bigcup Y$. Each vertex in the same subset ($X$ or $Y$) isn’t connected by an edge and only vertices in different subsets can be connected by an edge. We define the function Deduplicate and Count or $DC(\cdot)$ which, given a list of elements, counts the occurrences of each unique element in the list. It outputs a list of tuples of the form $(element,count)$ for each unique element in the list.
3.4. Construction Phase
@@
[Illustrated Training Data] Keyphrases Search Count Ranking audeze maxwell 1 audeze headphones 2 gaming headphones xbox 3 wireless headphones xbox 4 bluetooth wireless headphones 5   [Bipartite Graph derived from Illustrated Data] Refer to caption
Figure 3. Illustration of GraphEx’s construction phase. (i) a set of keyphrases with their search volume rank,(ii) shows the bipartite graph constructed from the set in (i).
An example of a constructed Bipartite Graph is shown in Figure 3. Each vertically stacked vertex belong to the same subset. The left set of vertices are the words/tokens and the right set are the keyphrases. The vertices are shown as strings here for presentation, but during implementation, the integer IDs are used. Each tokenized4 word is connected to the keyphrase that it is a part of. The graph is stored in Compressed Sparse Row (CSR) format, which occupies the least amount of space. Each word/token can be accessed in unit time whereas the adjacencies of a word can be traversed in $O(d)$ where $d$ is the degree of the word or the number of keyphrases that contain that word. A map type data-structure is used to associate the leaf category ID to the CSR structure for each graph.
The keyphrase’s Recall and Search Count [^5] Defined in section 3.3 are stored in separate arrays. So given a keyphrase ID $l$, $R(l)$ and $S(l)$ will directly index into the arrays and return the values taking unit time. The space occupied by each leaf category graph depends linearly on the number of unique words and edges, as CSR structure occupies $|X|+|E|$ space. The count of edges $|E|$ depends on the sum total of occurrence of each word in the keyphrases/labels which is difficult to generalize and depends on the datasets. Separate graphs for each leaf category help in recommending more relevant keyphrases which becomes more clear in the next section.
3.5. Inference Phase
Given a test item $T$ and a leaf category $l$ with the tokenized words in the title as $T={w_{1},w_{2},...,w_{t}}$, the goal is to extract a list of keyphrases in decreasing order of relevance to the item. GraphEx’s recommendation is based on permuting the word in the item’s title as discussed in section 1. To enable this, the Inference Phase is divided into two steps: Enumeration step that generates keyphrases from words of title and the Ranking step that ranks the keyphrases in order of relevance to the item.
3.5.1. Enumeration Step
GraphEx first determines the Bipartite Graph $G_{l}(V,E)$ that corresponds to the leaf category $l$ of the input item. The corresponding graph $G_{l}$ can be obtained in $O(1)$ time if a hashing data-structure is used for to map the leaf categories to the graphs defined in section 3.4. The step first tokenizes the item’s title into words and uses them as input along with the graph $G_{l}$ in the Algorithm 1. Lines 3-5 of the algorithm map the tokenized words of $T$ using the bipartite graph $G_{l}$ to the labels/keyphrases. Let’s look at an example to understand this process. Given an item “audeze maxwell gaming headphones for xbox”, we highlight the corresponding words on the left in the illustrated figure 3. The keyphrases ($l$) connected to the highlighted words are candidates for recommendation and are collected in $C_{L}$ in algorithm 1. Line 6 uses the $DC$ function to de-duplicate and count the redundancies in the candidate keyphrases. E.g. in figure 3 the keyphrase “audeze maxwell” is connected to two words “audeze” and “maxwell”, whereas “gaming headphones xbox” is connected to 3 words. Hence, after the execution of line 6, it results in the duplication count of 2,2,3,2, and 1 in the given order for each of the keyphrases on the right side of the illustrated figure 3. The count indicates the number of words in the keyphrase that are common with the item title $T$.
@@ 3.5.1. Enumeration Step
The next part of the Enumeration step generates a tuple corresponding to each label in $C_{L}$ using lines 7-9 in algorithm 1. We define the function Label Title Alignment or $LTA$ that uses the common word count (or duplication count) $c=|T\cap l|$ between the title $T$ and the label $l$ as $LTA(l,c)=\frac{c}{|l|-c+1}$. The LTA ratio is the second element of the tuple or the first attribute of the label $l$. The two attributes are the Search $S(l)$ and Recall count $R(l)$ of the label. The tuples generated by this process are returned in $C_{R}$. The time complexity of this step primarily depends on lines 3-5 due to restriction on prediction count which we discuss later in Section 3.6. The time complexity can be uncertain to determine due to the varying number of edges for each word. For simplification, we consider the average degree of each word as $d_{avg}=\frac{|E|}{|X|}$. Then asymptotically the time taken to gather the candidate labels for each word of the item title $T$ is $O(|T|.d_{avg})$. Modeling the problem as a Bipartite graph helps to efficiently permute all the words in the title $T$ while only generating permutations that are valid keyphrases.
3.5.2. Ranking Step
@@
In this step, the candidate labels in $C_{R}$ are sorted in the non-increasing order of the first attribute or second tuple element LTA and to break ties, $S(l)$ and subsequently $R(l)$ is used. While tie-breaking, those keyphrases are preferred that have higher search counts and lower recall counts. Higher search counts will have more clicks while lower recall count indicates the keyphrases have fewer items associated with them. So, when a keyphrase is input by a buyer, the search engine displays relatively fewer items, boosting click probability per item. The LTA function was designed to provide a higher score to those keyphrases that have less words in the label that aren’t part of the title. Let’s compare two keyphrases from the figure 3, “audeze maxwell” and “wireless headphones xbox”, both have 2 words in common with the sample title shown in section 3.5.1. The first’s LTA is $\frac{2}{1}$ and second’s is $\frac{2}{2}$, thus ranking “audeze maxwell” higher. LTA minimizes the risk involved by preferring those keyphrases that have more complete information (or more matching words).
3.6. Implementation Details
@@
The edges of the bipartite graph of each leaf category are constructed as tuples, sorted and then de-duplicated based on their IDs which are finally stored in the CSR format. The space complexity is linear in the number of edges for each graph given by $O(|X|\cdot d_{avg})$ where $X\in V$, which is the set of unique words in all the keyphrases for the leaf category. The words and the labels are represented as unsigned integers to occupy minimal space and convert string comparisons to integers ones. Therefore, comparing two words or two labels takes $O(1)$ time. The construction phase does not involve any weight updates or hyper-parameter training, making it quite fast and efficient.
@@
A drawback of directly using the algorithm 1 is the large number of keyphrases that are generated in the initial $C_{L}$. This results in a poly-logarithm time complexity for line 6 in the algorithm. To circumvent this we used count arrays to calculate the redundancies of each unit keyphrase. The space taken for the storage of $C_{L}$ and the count array is approximately $2|Q_{l}|$. A predetermined number of keyphrases (10-20) are generated for a given test instance during the inference phase. So, after the counting in line 6, the number of unique keyphrases in $C_{L}$ is pruned based on this requirement. This is done by first grouping each keyphrase with similar counts, then restricting the number of groups such that the sum of group sizes is equal to the required number of predictions. Groups with larger keyphrase redundancy counts are preferred, and all the keyphrases of the threshold group are included even if the group size overflows the number of required predictions. Thus the time complexity of the Enumeration step remains as $O(|T|.d_{avg})$. Though the sorting in the Ranking step seems expensive, the list length is always approximately a constant because of $|C_{L}|=|C_{R}|$. This is due to the restriction on prediction count as mentioned above thereby not contributing asymptotically to overall time complexity.
4. Experimentation and Results
@@
We perform experiments on representative datasets from eBay and compare our model’s results with representative models in production at eBay. We first describe our experimental setup, the datasets we use, and the models we compare in Subsection 4.1. Then, we analyze the results of each of our accuracy experiments in Subsection 4.2 and show the execution performance of each model in  4.3. Next, we describe the deployment in production in section 4.4 and its impact in section 4.5.
4.0.1. Setup and Datasets
GraphEx is implemented for multi-core systems without requiring any GPUs. Its inference part is implemented on C++ ($\geq$ g++-9.3.0) using OpenMP threading with Python wrappers using pybind11. The construction part is implemented in Python ($\geq$ 3.7); due to its lightweight approach since the construction doesn’t require large resources and takes much less time. We used a system with 4 Intel Xeon Gold 6230R CPUs with 2 sockets each containing 20@2.10GHz cores, and 500 GB of RAM for the analysis. GraphEx employs coarse-grained multi-threading, assigning each input’s inference to an individual thread. We launched 20 threads with compact pinning to occupy only a single socket sufficient for our dataset size.
@@ 4.0.1. Setup and Datasets
Table 1. Details of three representative categories of eBay.
4.1. Experimentation Details
We contrast our model against notable models used at eBay for keyword recommendation, as elaborated in section 2. We present findings on three product meta-categories from eBay, each symbolizing a classification of large, medium, and small categories. The classification is determined by the count of items and the quantity of unique keyphrases within each meta-category. Table 1 shows the anonymized categories and their details. Even though our methodology does not require knowledge of the items or their meta-data, the XMC models require them, hence we show their numbers for perspective. Our data curation and analysis are limited to eBay, due to the absence of any publicly available keyword recommendation datasets from e-commerce advertisement platforms.
@@ 4.1. Experimentation Details
Figure 4. The average counts of relevant head/tail and irrelevant keyphrases per item are shown for each model.
@@
The data is collected from search logs for the duration of one year for both XMC models and GraphEx. For XMC, the item-keyphrase pairs are constrained based on their co-occurrence count, number of buyer clicks/purchases, etc. The curated unique keyphrase count shown in the third column of table 1 contains both the head and tail keyphrases and is incorporated by XMC models. On the other hand, GraphEx’s data curation for training, aggregates keyphrases without looking at any association with the items. It restricts the keyphrases [^6] Shown in the right most column of table 1. to contain a higher number of head and a lower number of tail keyphrases using the curation process described in Section 3.2. Generally, keyphrases that on an average weren’t searched atleast once per day were filtered for GraphEx [^7] This restriction was relaxed for CAT_3 that didn’t have sufficient keyphrases.. For testing, we sampled a set of 1000, 400 and 200 items from actively listed items on eBay.com for the categories CAT_1, CAT_2, and CAT_3 respectively. We also computed the search count of each unique keyphrase by considering a 15 day duration different from the one year duration for the training set. This removes any bias that models have based on their training data. For each of the test items, all the models generate a variable number of keyphrases with a limit of 40.
4.1.1. Traditional Metrics
Typically, metrics like Precision, Recall, F1 and so on are used for comparing recommendation models. These metrics facilitate comparison by emphasizing retrieval capability, which is suitable for XMC tagging models, but does not pertain to extraction or generation models. There are three main issues with using these metrics: Prediction Diversity, Ground Truth Incompleteness/Uncertainty, and Ground Truth Bias. Let us take an example instance $T$, associated with a ground truth label $k1$ in the training set. All models make certain choices to increase the probability of predicting $k1$ for inputs that are similar to $T$. This aligns the tagging-based models to predict a similar subset of labels for $T$, thus reducing diversity among the predictions of different models. Another facet is the assumption that the curated ground-truth labels for each instance or data point are correct. Labels for some instances or items might be missing due to rigid thresholds in the curation process. Hence, a metric that measures retrieval ability hinders the performance of unsupervised extraction or generation tasks that don’t rely on ground truth labels for items. Finally, if $k1$ is a biased label for $T$, tagging models tend to replicate the bias in their predictions. Improving diversity and reducing biases is possible through the alteration of techniques, though this would complicate their implementation and the process of comparison.

### v3

@@ III-C Terms and Notations
We consider a set of unique keyphrases termed as $Q={k_{1},k_{2},...,k_{K}}$. Each keyphrase $k_{i}$ can be considered as a set of words ${w_{1},w_{2},...,w_{l}}$, where $w_{*}$ are tokenized [^3] The tokenization scheme can be anything as long as string comparison functions are well-defined and consistent for that scheme. By default we consider space-delimited tokenization. from the keyphrase string $k_{i}$. Each $k_{i}$ is further associated with a Leaf category $l$, Recall Count or Rank $R$ and Search Count or Rank $S$. Given a test item’s title $T$, the goal is to recommend a subset of keyphrases from $Q$ that are relevant to $T$. We can consider the title as a string with tokenized3 words $T={w_{1},w_{2},...,w_{t}}$ similar to a keyphrase, but titles are generally longer than the keyphrases. We denote a graph $G(V,E)$ where $V$ is the set of vertices and $E$ is the set of edges. Each edge $e\in E$ is denoted by a pair of vertices $e=(v_{1},v_{2})$ indicating a connection between the vertex pair. In a Bipartite Graph, the set of vertices $V$ are divided into a pair of disjoint subsets $V=X\bigcup Y$. Each vertex in the same subset ($X$ or $Y$) isn’t connected by an edge and only vertices in different subsets can be connected by an edge. We define the function Deduplicate and Count or $DC(\cdot)$ which, given a list of elements, counts the occurrences of each unique element in the list. It outputs a list of tuples of the form $(element,count)$ for each unique element in the list.
III-D Construction Phase
@@
(a) Illustrated Training Data
(b) Bipartite Graph derived from Illustrated Data
Figure 3: Illustration of GraphEx’s construction phase. (i) a set of keyphrases with their search volume rank, (ii) shows the bipartite graph constructed from the set in (i).
An example of a constructed Bipartite Graph is shown in Figure 3. Each vertically stacked vertex belong to the same subset. The left set of vertices are the words/tokens and the right set are the keyphrases. The vertices are shown as strings here for presentation. Each tokenized3 word is connected to the keyphrase that it is a part of. The graph is stored in Compressed Sparse Row (CSR) format, which occupies the least amount of space. Each word/token can be accessed in constant time whereas the adjacencies of a word can be traversed in $O(d)$ where $d$ is the degree of the word or the number of keyphrases that contain that word. A map type data-structure is used to associate the leaf category ID to the CSR structure for each graph.
@@
The keyphrase’s Recall and Search Count [^4] Defined in section III-C are stored in separate arrays. So given a keyphrase ID $l$, $R(l)$ and $S(l)$ will directly index into the arrays and return the values taking unit time. The space occupied by each leaf category graph depends linearly on the number of unique words and edges, as CSR structure occupies $|X|+|E|$ space. The count of edges $|E|$ depends on the sum total of occurrence of each word in the keyphrases/labels which is difficult to generalize and depends on the datasets. Separate graphs for each leaf category help in recommending more relevant keyphrases which becomes more clear in the next section.
III-E Inference Phase
Given a test item $T$ and a leaf category $l$ with the tokenized words in the title as $T={w_{1},w_{2},...,w_{t}}$, the goal is to extract a list of keyphrases in decreasing order of relevance to the item. GraphEx’s recommendation is based on permuting the words in the item’s title as discussed in section I. To enable this, the Inference Phase is divided into two steps: Enumeration step that generates keyphrases from words of title and the Ranking step that ranks the keyphrases in order of relevance to the item.
III-E1 Enumeration Step
GraphEx first determines the Bipartite Graph $G_{l}(V,E)$ that corresponds to the leaf category $l$ of the input item. The corresponding graph $G_{l}$ can be obtained in $O(1)$ time if a hashing data-structure is used to map the leaf categories to the graphs defined in section III-D. The step first tokenizes the item’s title into words and uses them as input along with the graph $G_{l}$ in the Algorithm 1. Lines 3-5 of the algorithm map the tokenized words of $T$ using the bipartite graph $G_{l}$ to the labels/keyphrases. Let’s look at an example to understand this process. Given an item “audeze maxwell gaming headphones for xbox”, we highlight the corresponding words on the left in the illustrated Figure 3(b). The keyphrases ($l$) connected to the highlighted words are candidates for recommendation and are collected in $C_{L}$ in Algorithm 1. Line 6 uses the $DC$ function to de-duplicate and count the redundancies in the candidate keyphrases. E.g. in Figure 3(b) the keyphrase “audeze maxwell” is connected to two words “audeze” and “maxwell”, whereas “gaming headphones xbox” is connected to 3 words. Hence, after the execution of line 6, it results in the duplication count of 2,2,3,2, and 1 in the given order for each of the keyphrases on the right side of the illustrated Figure 3(b). The count indicates the number of words in the keyphrase that are common with the item title $T$.
@@ III-E1 Enumeration Step
The next part of the Enumeration step generates a tuple corresponding to each label in $C_{L}$ using lines 7-9 in Algorithm 1. We define the function Label Title Alignment or $LTA$ that uses the common word count (or duplication count) $c=|T\cap l|$ between the title $T$ and the label $l$ as $LTA(l,c)=\frac{c}{|l|-c+1}$. The LTA ratio is the second element of the tuple or the first attribute of the label $l$. The two attributes are the Search $S(l)$ and Recall count $R(l)$ of the label. The tuples generated by this process are returned in $C_{R}$. The time complexity of this step primarily depends on lines 3-5 due to restriction on prediction count which we discuss later in Section III-F. The time complexity can be uncertain to determine due to the varying number of edges for each word. For simplification, we consider the average degree of each word as $d_{avg}=\frac{|E|}{|X|}$. Then asymptotically the time taken to gather the candidate labels for each word of the item title $T$ is $O(|T|.d_{avg})$. Modeling the problem as a Bipartite graph helps to efficiently permute all the words in the title $T$ while only generating permutations that are valid keyphrases.
III-E2 Ranking Step
@@
In this step, the candidate labels in $C_{R}$ are sorted in the non-increasing order of the first attribute or second tuple element LTA and to break ties, $S(l)$ and subsequently $R(l)$ is used. While tie-breaking, those keyphrases are preferred that have higher search counts and lower recall counts. Higher search counts will have more clicks while lower recall count indicates the keyphrases have fewer items associated with them. So, when a keyphrase is input by a buyer, the search engine displays relatively fewer items, boosting click probability per item. The LTA function was designed to provide a higher score to those keyphrases that have less words in the label that aren’t part of the title. Let’s compare two keyphrases from the Figure 3(b), “audeze maxwell” and “wireless headphones xbox”, both have 2 words in common with the sample title shown in section III-E1. The first’s LTA is $\frac{2}{1}$ and second’s is $\frac{2}{2}$, thus ranking “audeze maxwell” higher. LTA minimizes the risk involved by preferring those keyphrases that have more complete information (or more matching words).
III-F Implementation Details
@@
The edges of the bipartite graph of each leaf category are constructed as tuples, sorted and then de-duplicated based on their IDs which are finally stored in the CSR format. The space complexity is linear in the number of edges for each graph given by $O(|X|\cdot d_{avg})$ where $X\in V$, which is the set of unique words in all the keyphrases for the leaf category. The words and the labels are represented as unsigned integers to occupy minimal space and convert string comparisons to integer ones. Therefore, comparing two words or two labels takes $O(1)$ time. The construction phase does not involve any weight updates or hyper-parameter training, making it quite fast and efficient.
Generally, the leaf category ids within a meta category are unique. There are a large number of leaf categories ($>100000$ in total). Models like fastText and Graphite were trained per meta category which are approximately 120. Other non-rule based models like SL were even more granular with item co-click based associations. GraphEx handles the subcategories internally and doesn’t need to train one model per leaf category, minimizing computational costs. Also, if the leaf categories are unique across various meta categories then only model will be required for the entire group. The leaf categories help GraphEx in recommending relevant keyphrases as generally the items and keyphrases in the leaf categories belong to the same product.
@@
A drawback of directly using the Algorithm 1 is the large number of keyphrases that are generated in the initial $C_{L}$. This results in a poly-logarithm time complexity for line 6 in the algorithm. To circumvent this we used count arrays to calculate the redundancies of each unit keyphrase. The space taken for the storage of $C_{L}$ and the count array is approximately $2|Q_{l}|$. A predetermined number of keyphrases (10-20) are generated for a given test instance during the inference phase. So, after the counting in line 6, the number of unique keyphrases in $C_{L}$ is pruned based on this requirement. This is done by first grouping each keyphrase with similar counts, then restricting the number of groups so that the sum of group sizes is equal to the required number of predictions. Groups with larger number of keyphrase redundancy counts are preferred, and all keyphrases in the threshold group are included even if the group size exceeds the number of required predictions. Thus, the time complexity of the Enumeration step remains as $O(|T|.d_{avg})$. Although the sorting in the Ranking step seems expensive, the list length is always approximately a constant due to $|C_{L}|=|C_{R}|$. This is due to the restriction on prediction count as mentioned above thereby not contributing asymptotically to overall time complexity.
III-G Interpretability
The applications in E-Commerce domain frequently require that a model be interpretable. This helps to comprehend the rational behind its predictions and decision process. In our use case, it is essential to trace where the words in the keyphrases come from. Neural network models typically require converting input text into vectors, which often obscures the contribution of individual tokens to the decisions. Interpretability techniques such as LIME and SHAP offer post hoc explanations, treating a Deep Neural Network as a black-box. They also require much effort to figure out the contributions of each input feature.
Unlike black-box models, the GraphEx algorithm has 3 transparent phases: keyphrase curation, keyphrase mapping, and ranking. The data curation process gives perspective as to how the keyphrases in GraphEx’s label set were curated. The keyphrase mapping phase details how GraphEx’s candidate keyphrases were mapped from the keyphrases extracted from the item’s title to GraphEx’s candidate keyphrases. The ranking algorithm which then ranks the mapped candidates is transparent as well. It uses Label Title Alignment (LTA) outlined in Section III-E1 which is a token-based algorithm ensuring that the majority of tokens in the keyphrases match the title. This ensures that GraphEx’s predicted keyphrases are explainable and interpretable.
IV Experimentation and Results
@@
We perform experiments on representative datasets from eBay and compare the results of our model with the described models (Section II) in production at eBay. We first describe our experimental setup, the datasets we use and the models we compare in Section IV-B. Next, we describe our evaluation framework in Section IV-C on how we determine relevance and the metrics using relevance for performance comparison. Then, we analyze the results of each model’s performance in sections IV-D and IV-E, ablation studies in IV-F and the execution performance of each model in  IV-G. Finally, we describe the deployment in production in Section IV-H and its impact in section IV-I.
@@
TABLE II: Details of three representative meta categories of eBay.
@@
Figure 4: The average counts of relevant head/tail and irrelevant keyphrases per item are shown for each model.
IV-A Setup and Datasets
GraphEx is implemented for multi-core systems without requiring any GPUs. Its inference part is implemented on C++ ($\geq$ g++-9.3.0) using OpenMP threading with Python wrappers using pybind11. The construction part is implemented in Python ($\geq$ 3.7); due to its lightweight approach since the construction does not require large resources and takes much less time. We used a system with 4 Intel Xeon Gold 6230R CPUs with 2 sockets each containing 20@2.10GHz cores, and 500 GB of RAM for the analysis. GraphEx employs coarse-grained multithreading, assigning each input’s inference to an individual thread. We launched 20 threads with compact pinning to occupy only a single socket sufficient for our dataset size.
IV-B Experimentation Details
We present findings on three product meta categories from eBay, each symbolizing a classification of large, medium, and small categories. The classification is determined by the count of items and the quantity of unique keyphrases within each meta-category. Table II shows the anonymized categories and their details. Even though our methodology does not require knowledge of the items or their meta-data, the XMC models require them, and hence we show their numbers for perspective. Our data curation and analysis are limited to eBay, due to the absence of any publicly available keyphrase recommendation datasets from e-commerce advertisement platforms.
@@ IV-B Experimentation Details
The data is collected from search logs for the duration of one year for XMC models and 6 months for GraphEx. For XMC, the item-keyphrase pairs are constrained based on their co-occurrence count, number of buyer clicks/purchases, etc. The curated unique keyphrase count shown in the third column of table II contains both the head and tail keyphrases and is incorporated by XMC models. On the other hand, GraphEx’s data curation for training, aggregates keyphrases without looking at any (clicked-based) association with the items. It restricts the keyphrases [^5] Shown in the right most column of table II. to contain a higher number of head and a lower number of tail keyphrases using the curation process described in Section III-B. Generally, keyphrases that on an average weren’t searched at least once per day were filtered for GraphEx [^6] This restriction was relaxed for CAT_3 that didn’t have sufficient keyphrases..
For testing, we sampled a set of 1000, 400 and 200 items from actively listed items on eBay.com for the categories CAT_1, CAT_2, and CAT_3 respectively. We also computed the search count of each unique keyphrase by considering a 15 day duration different from the one year duration for the training set. This removes any bias that models have based on their training data. For each of the test items, all the models generate a variable number of keyphrases with a limit of 40.
IV-C Evaluation Framework
We described here our evaluation framework that tackles the challenges in Section I-A3. Summarizing the two major problems are 1) Lack of ground truths due to sparse data and MNAR-led biases in click data and 2) Model Convergence due to similar training data and evaluation. Ideally, metrics should compare the relevancy of the predictions to the input text without limiting the comparison to a set of predefined labels/keyphrases. However, it is difficult to determine the relevance of predictions without any prior labels or the absence of negative labels [^7] Absence of click does not imply irrelevance, see Section I-A2.